A transmission line carries a current of 100A. What would be the magnetic field B at a point on the road if the wire is 8m above the road?

Grade:10

1 Answers

Khimraj
5 years ago
B = uoI/2\pi a = (uo/4\pi)(2I/a) = 10-7*2*100/8 = 2.5\mu T
Hope it clears. If you like answer then please approve it.